OpenStudy (anonymous):

Which of the following equations describes the line shown below? Check all that apply. (-5.-6) (1,6)

OpenStudy (anonymous):

A. y=2x+2 B. y=2x+4 C. y-6=2(x-1) D. y-6=2(x-5) E. y+6=2(x+5) F. y+6=2(x+1)

OpenStudy (anonymous):

@jcpd910 Help

OpenStudy (anonymous):

I can help in about twenty minutes k?

OpenStudy (anonymous):

k

OpenStudy (anonymous):

@jcpd910

OpenStudy (anonymous):

Mk let's go

OpenStudy (anonymous):

Remember, these are your two points right? (-5,-6) (1,6)

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

And the points are written like this: (x,y) remember?

OpenStudy (anonymous):

kinda

OpenStudy (anonymous):

Okay, so you can do two things: Replace x with -5 in every equation and see if it comes out true Or you can convert this into an equation after calculating the slope

OpenStudy (anonymous):

Calculating the slope is much faster.

OpenStudy (anonymous):

So, which do you want to do?

OpenStudy (anonymous):

calculating the slope

OpenStudy (anonymous):

Okay, so the slope formula is: \[m = \frac{ y _{1} - y _{2} }{ x _{1}-x _{2} }\] Based on the points you are given: (-5,-6) (1,6) \[x _{1} = -5, x _{2} = 1, y _{1} = -6, y _{2} = 6\]

OpenStudy (anonymous):

So plug in those values.

OpenStudy (anonymous):

ummmm

OpenStudy (anonymous):

Tell me what y1 is, I just wrote it out.

OpenStudy (anonymous):

y=-6

OpenStudy (anonymous):

y1 = -6, we have two y values so we put y1 and y2 so we know which is which.

OpenStudy (anonymous):

So we sub the values in like this: \[m = \frac{ y _{1} - y _{2} }{ x _{1}-x _{2} }\] \[x _{1} = -5, x _{2} = 1, y _{1} = -6, y _{2} = 6\] \[m = \frac{ -6 - 6 }{ -5-1 }\]

OpenStudy (anonymous):

See?

OpenStudy (anonymous):

so what would the answer be

OpenStudy (anonymous):

Well what's -6 -6 and what is -5 - 1?

OpenStudy (anonymous):

???

OpenStudy (anonymous):

x1=-5 x2=1 y1=-6 y2=6

OpenStudy (anonymous):

Great, but what is: -6 - 6 = -5 - 1 =

OpenStudy (anonymous):

Eric?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

Okay, what is 0 - 6

OpenStudy (anonymous):

0-6=-6

OpenStudy (anonymous):

So what is -6 - 6?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

no...

OpenStudy (anonymous):

jus playin -12

OpenStudy (anonymous):

lol k

OpenStudy (anonymous):

so what is -5 - 1?

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

Yup! okay, so now we have: \[m = \frac{ -12 }{ -6 }\]

OpenStudy (anonymous):

Remember, when you divide a negative by a negative, the result is positive. So if -4 / -2 = 2, what does -12 / -6 =?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

Now, what is 12 / 6?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

so -12 / -6?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

right, so we have our slope. \[m = 2\]

OpenStudy (anonymous):

Now we can put that into point-slope form: y = mx + b

OpenStudy (anonymous):

y = 2x + b

OpenStudy (anonymous):

but we need to solve for b now.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

So we use the points we're given: (-5.-6) (1,6) And the slope, 2, to find our y intercept. x = -5 y = -6 x = -4 y = -4 x = -3 y = -2 x = -2 y = 0 x = -1 y = 2 x = 0 y = 4 x = 1 y = 6

OpenStudy (anonymous):

When x = 0, you have your y intercept. Which one is our y intercept for this line?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

x = 0, y = ???

OpenStudy (anonymous):

4

OpenStudy (anonymous):

right, so let me draw it out so you understand.

OpenStudy (anonymous):



OpenStudy (anonymous):

so now we have: y = mx + b y = 2x + 4

OpenStudy (anonymous):

yeaa

OpenStudy (anonymous):

Which automatically means b is right, but we can check the others.

OpenStudy (anonymous):

y-6=2(x-1) y = 2(x - 1) + 6 We already know this is wrong, because b = the y intercept, and right now b = 6. Our y intercept is 4.

OpenStudy (anonymous):

y - 6 = 2(x-5) y = 2(x - 5) + 6 Wrong

OpenStudy (anonymous):

And the last two will have -6 as slopes, which is still wrong. The only correct answer is B. :D

OpenStudy (anonymous):

thank you so much 4 da help @jcpd910